### v1

# AI-Based Beam-Level and Cell-Level Mobility Management for High Speed Railway Communications

Abstract
High-speed railway (HSR) communications are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services. The high speed of trains creates rapidly time-varying wireless channels, increases the signaling overhead, and reduces the system throughput, making it difficult to meet the growing and stringent needs of HSR applications. In this article, we explore artificial intelligence (AI)-based beam-level and cell-level mobility management suitable for HSR communications, including the use cases, inputs, outputs, and key performance indicators (KPI)s of AI models. Particularly, in comparison to traditional down-sampling spatial beam measurements, we show that the compressed spatial multi-beam measurements via compressive sensing lead to improved spatial-temporal beam prediction. Moreover, we demonstrate the performance gains of AI-assisted cell handover over traditional mobile handover mechanisms. In addition, we observe that the proposed approaches to reduce the measurement overhead achieve comparable radio link failure performance with the traditional approach that requires all the beam measurements of all cells, while the former methods can save 50% beam measurement overhead.
Index Terms:
Artificial intelligence, beam-level mobility management, cell-level mobility management.I Introduction
High speed railway (HSR) provides people with safe, fast, comfortable and economical modes of transportation, and has been widely developed around the world. Nevertheless, mobile communication systems are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services. Emerging applications, such as railway multimedia dispatching, high-definition video monitoring on board and trackside, have posed stricter requirements on the reliability, data rate, and spectrum efficiency of the next-generation, dedicated mobile communication systems for HSR. However, the high speed of trains entails formidable challenges due to the time-varying nature of wireless channels in HSR communications. These challenges include the increased signaling overhead and reduced system throughput.
Thanks to the abundant spectrum available at higher frequencies, communication systems using the millimeter wave (mmWave) technology avail of faster data transmission speeds and reduced latency compared to traditional low-frequency communications. For HSR communications, mmWave technology underpins increased data transmission rates that meet the demands of onboard applications and services, and provides low latency that is indispensable for real-time communication requirements. However, mmWave communication systems suffer from significant path loss, which restricts the effective range of transmission. To tackle this problem, beamforming with large-scale antenna arrays can be employed to generate finely directional beams, mitigating propagation loss and expanding signal coverage. However, the high mobility of trains entails pressing challenges in mobility management for HSR communications. While being connected, a train has to address two types of mobility, i.e., the cell-level and beam-level mobility as specified in the 3rd Generation Partnership Project (3GPP) Release 17 for 5G systems[1]. Beam-level mobility involves dynamically switching beams in which the train establishes a link path to a new beam within the same cell, ensuring communication with high performance. On the other hand, cell-level mobility focuses on cell handover between different cells. As trains move from one coverage zone to another, it is crucial to have a smooth cell handover to sustain continuous connectivity experience.
The management of mobility at the beam and cell level in 5G is based on the use of periodic channel state information reference signals (CSI-RS) and synchronization signal blocks (SSBs). Unfortunately, the resulting large latency and high overhead are prohibitive for HSR communications, and, hence, innovative approaches are urgently needed to address these limitations. Deep neural networks (DNNs) have found extensive applications in areas, such as computer vision and natural language processing, owing to their exceptional ability to extract features from intricate and high-dimensional data. In recent years, DNNs have also proven their effectiveness in capturing the complex and nonlinear characteristics of wireless channels to facilitate beam-level mobility management. For example, a multi-layer long short term memory (LSTM) network was utilized to predict the optimal beam change for future instances [2]. Studies on advanced artificial intelligence (AI)-based cell-level mobility management are relatively few. For example, [3] utilized convolutional neural network (CNN)-based auto-encoders for dimension reduction and LSTM networks to predict the appropriate base station (BS) for cell handover. Recently, the 3GPP 5G NR organization has launched studies on AI-based cell-level mobility management for highly reliable handovers in Release 19. However, these AI-based mobility management methods fail to exploit the specific characteristics in HSR communications. For instance, trains typically maintain stable high-speed motion on fixed tracks and have different speed levels to meet different needs. Mobility management methods for HSR communications are expected to leverage case-specific characteristics in the design phase or, alternatively, learn from data to yield improved performance in a data-driven manner.
In this article, we present the latest progress in AI-based beam-level and cell-level mobility management suitable for HSR communications. In Section II, we reveal the potential of AI-based spatial-temporal beam prediction for HSR communications. Particularly, in comparison to the traditional down-sampling spatial beam measurements, we show that the compressed spatial multi-beam measurements via compressive sensing (CS) can lead to improved spatial-temporal beam prediction in HSR communications. In Section III, we investigate AI-based cell-level mobility management for HSR scenarios, including use cases, the input/output of the AI model, evaluation indicators, and enhanced schemes. Particularly, we demonstrate the performance gains of AI-assisted cell handover over traditional cell handover mechanisms for HSR scenarios. Section V introduces the challenges and opportunities encountered in AI-based mobility management, followed by our conclusions in Section VI.
II AI-Based Beam-Level Mobility Management for HSR Communications
II-A Use Cases in HSR Communications
Beam-level mobility enables a user equipment (UE) to communicate with a BS via the optimal beam pair, which involves low-level beam operations rather than handover. In HSR communications, with some coarse-grained beam quality measurements of the serving cell, AI can be utilized to estimate the finely-grained optimal beam pair in the current instance or predict the optimal beam pair in future instances, which leads to reduced measurement overhead and enhanced throughput.
There are two primary use cases for AI models applied in beam-level mobility management, i.e., temporal beam prediction and spatial beam prediction. For temporal beam prediction, where the beam quality information of previous instances is utilized to predict the quality of beams in one or more future instances, most AI-based methods employ recurrent neural networks (RNNs) and their variants, such as LSTM and gated recurrent unit (GRU) models. This is driven by the ability of RNN models to identify temporal patterns in time series data, allowing them to gather extensive information. A two-layer convolutional LSTM network can be used for predicting the beam quality across spatial and temporal dimensions as detailed in [4]. The combination of a CNN and LSTM network was proposed in [5] for predicting the optimal beam ID across temporal dimensions. A multi-layer LSTM network was applied in [6] for forecasting the optimal future beam. Spatial beam prediction exploits spatial correlations among adjacent beams to estimate the quality of all beams using incomplete beam data. CNNs and networks with fully connected (FC) layers have been commonly employed for this purpose [7, 8, 9]. Due to the high mobility in HSR communications, the extensive signaling overhead and the time required for frequent beam sweeping become prohibitive. For these reasons, temporal beam prediction offers greater benefits in terms of resource conservation and stable performance in HSR communications.
II-B Inputs and Outputs in HSR Communications
Different tasks in beam-level mobility management require distinct input/output into/of the AI models in HSR communications. According to the ongoing discussions within 3GPP, there are two vital concepts for beam management, namely Set A and Set B. Set A comprises all narrow beams from the predefined codebook, while Set B encompasses the beams that have been measured and are utilized as inputs into the AI models. Various alternative methods exist with respect to the relationship between Set A and Set B. For instance, the AI model could receive the input of reference signal received power (RSRP) measurements of partial narrow beams, making Set B as a subset of Set A. This is the most commonly used method in HSR communications, which greatly reduces the measurement overhead. Another option is to utilize RSRP measurements of wide beams to estimate the quality of narrow beams, where Set B and Set A are distinct. In addition, Set B can be the same as Set A, whenever the RSRP measurements of all narrow beams are the inputs into the AI models.
Note that the first two types of inputs are employed in temporal and spatial beam prediction models, while the last is only for temporal beam prediction. Instead of using RSRP, the complex received signal can be used as input into the AI model, which will be discussed in the CS-based method in the next subsection. In addition, supplementary auxiliary information can be incorporated to improve the prediction accuracy for HSR communications. In [10], the position of the UE, its mobility vector, the signal-to-noise ratio (SNR), and the beam index currently in use were utilized to forecast the state of the link connection. Furthermore, measurements from light detection and range (LIDAR) and various sensors, images of the environment taken by cameras, and information from low-frequency channels could also be employed as extra inputs into the AI model to enhance the prediction accuracy.
There are two types of outputs in the AI model design for beam-level mobility management. The most common outputs are weights associated with different beams, where a higher weight suggests a better beam. Additionally, line-of-sight (LoS) is the primary mode of signal propagation in HSR communications and beam-selection is easier in the LoS setting. Therefore, the AI model can be utilized to output the estimated state of each path, i.e., LoS or non-LoS, to enhance the beam prediction [11, 9].
II-C AI Enhancement for HSR Communications
Beam-level mobility management enhancement involves three types of beam search methods: traditional methods including exhaustive search and hierarchical search, advanced numerical methods, and AI-based methods. The exhaustive beam search method examines all the candidate beams in the predefined codebook one by one to obtain beam quality information. The hierarchical beam search method provides a more efficient alternative, which first searches within a wide beam range and then further searches for a narrow beam within the angular range of the selected wide beam. However, traditional beam search algorithms suffer from a large training overhead, which is not suitable for HSR communications which require frequent beam switching. In addition, enhancement techniques, based on numerical methods have been proposed. More specifically, [12] exploited the tree search method for efficient beam training based on the hierarchical codebook, whilst [13] proposed an optimal beam training model based on Markov’s decision process. These numerical approaches mainly target general communication environments, making it challenging to tailor them for particular communication scenarios.
In HSR communications, every train follows a predetermined path and performs the same routes periodically. Given sufficient training data, AI models are expected to perform exceptionally well in managing beam-level mobility owing to this intrinsic regularity. For instance, AI-based super-resolution techniques can be employed for spatial beam prediction to reduce the measurement overhead in HSR communications. Neural networks can receive inputs in the form of low-resolution beam measurements, consisting of partial narrow beams or wide beams, and output high-resolution spatial narrow beam predictions. The super-resolution-based spatial beam prediction can be implemented using a sub-pixel convolutional layer paired with an up-sampling layer [4], or a combination of CNN and FC [7].
Nevertheless, these existing beam prediction solutions rely on the premise that the assessed beams are selected from a predefined subset of all the beams, akin to the down-sampling process. One limitation of beam measurement through down-sampling is that it does not capture any information from the beams that are not sampled, leading to inevitable performance loss. Rather than discarding certain beams outright, a potential approach to reduce this performance loss is to use CS that projects the channel into a lower dimensional space. Given some CS-based beam measurements, deep learning (DL) has been exploited for spatial beam prediction [14].
Here, we investigate DL-based spatial-temporal beam prediction using compressed multi-beam channel measurements. DL is employed to output the beam quality in the future slots using historical multi-beam channel measurements in the compressed domain, where the CS measurement matrix can be learned jointly. In the training phase, the complete beam channel information from historical instances serves as input into the model. The CS matrix serves as the first layer, namely the linear compression layer. The compressed channel information is fed into the remaining neural network to reconstruct the complete RSRP information of all candidate beams for a future instance. As a learnable parameter, the CS measurement matrix is jointly optimized during the training of the model. In the inference phase, the trained measurement matrices are applied at the transceivers and the compressed measurements are fed into the trained neural network to conduct beam prediction. Compared with the traditional sweeping method, the multi-beam compression offers to the beam prediction network richer information with the same number of beam measurements. The structure of the proposed AI model for beam-level mobility management, namely CSAI, is illustrated in Fig. 2(d).
Now, we compare the performance of AI-based spatial-temporal beam prediction methods, whose structures are shown in in Fig. 2. The AI-based methods in Fig. 2(a) [6] and Fig. 2(b) [7] employ the traditional LSTM and CNN architecture, respectively, to predict the high-resolution beams. Compared with the CSAI in Fig. 2(d), the AI-based method in Fig. 2(c) [4] has the same network structure except for the beam measurement module. In the experiments, we consider the UMa scenario for outdoor 30 GHz mmWave communications, while the UE follows a straight-line trajectory which aligns with the HSR scenario. Four distinct datasets, representing UE speeds of 60km/h, 120km/h, 350km/h and 500km/h, are utilized in the evaluation. These datasets comprise 40,000 training samples, 5,000 validation samples, and 5,000 test samples. The batch size is set to 256, and the training runs for a maximum of 800 epochs. The learning rate changes dynamically from to , utilizing the adaptive moment estimation optimizer. We consider the prediction accuracy for the best beam with a compression ratio111The compression ratio is defined as the ratio of the number of input beam measurements to the number of all beam measurements. of . The simple non-AI approach straightforwardly uses the best one of the partially measured present beams as the prediction. As shown in Fig. 3, the AI-based spatial-temporal beam prediction methods outperform significantly the traditional non-AI method. As the speed of the movement increases, all methods generally show an anticipated decrease in the prediction accuracy. The CSAI demonstrates superior performance for all the cases. At the speed of 500km/h, the CSAI is still able to yield a prediction accuracy of 83.76%.
III AI-Based Cell-level Mobility Management in HSR Communications
III-A Use Cases in HSR Communications
Cell-level mobility management deals primarily with cell handovers, initiating the transition of a user’s session to a new cell when the signal strength from the serving cell falls below a predefined threshold. Given the rapid movement in HSR communications, cell handovers occur frequently, and failures in these handovers would unavoidable compromise the user experience. As specified in 3GPP, the cell handover procedure relies on concrete criteria, such as event A3, which will be triggered when the difference between the layer 3 RSRP (L3-RSRP)222L3-RSRP represents the cell-level RSRP, obtained by averaging or maximizing the RSRP of all beam pairs in each cell. measurement of the target cell and the L3-RSRP measurement of the source cell exceeds a predefined offset. If the above criteria are met during the time to trigger (TTT) duration, the UE reports the measurements and the source cell initiates the handover preparation phase. Afterwards, the cell handover procedure initiates the execution after the UE receives the handover command message, whose task is to provide radio resource control (RRC) reconfiguration message to the target cell. The specific cell handover process is shown in Fig. 4. During state 1, the UE must sweep across all beams to acquire the L3-RSRP of adjacent cells for the TTT duration, leading to signal quality deterioration in the service cell, which are even more pronounced in HSR scenarios.
To maintain a stable communication quality for HSR communications, the most common use case for AI-based cell-level mobility management is to predict the RSRP of all adjacent cells in future TTT instances. The AI-based cell handover process is illustrated in Fig. 4, where option 1 and option 2 represent different deployment methods of AI models. In this case, one option for the source cell is to immediately execute a cell handover based on the AI model’s predicted target cell. Alternatively, the source cell can monitor the signal quality of both the target and source cells over a period of time and initiate the handover if some handover conditions are still met. This approach can effectively reduce the ping-pong handovers and enhance the communication stability, while it requires more handover time for signal detection. In HSR scenarios, where trains move in straight lines and the probability of ping-pong handovers is low, the delay introduced by extended handover time, resulting in significant communication quality degradation, becomes intolerable. Therefore, the first approach is better suited for cell-level mobility management in HSR communications.
III-B Inputs, Outputs, and KPIs in HSR Communications
For the HSR scenario, we explore various model inputs and outputs tailored to its specific requirements. One input option for AI models is the L3-RSRP of each cell at historical instances, with the output being either the L3-RSRP of each cell or the L3-RSRP of the top cells at future TTT instances. The dimension of L3-RSRP information is lower than that of L1-RSRP. However, acquiring L3-RSRP, which requires averaging data over multiple time slots, introduces extra latency which is not desired in HSR communications. Another viable approach involves using layer 1 RSRP (L1-RSRP)333L1-RSRP represents the beam-level RSRP, which is the RSRP of each beam pair of each cell. as both the input and the output of the AI models. The AI model output is then filtered to derive the L3-RSRP, which is used to execute cell handovers. For this option, the output L1-RSRP of AI model can be applied to both beam-level and cell-level mobility management. Nevertheless, L1-RSRP encompasses extensive information that exceeds the requirements for handover, while the correlation between the input and output L1-RSRP is intricate, thereby amplifying the learning challenge for the AI models. Moreover, leveraging L1-RSRP to forecast L3-RSRP via AI can be an advantageous strategy. This method reaps the benefits of the above two approaches, making use of the abundant data in L1-RSRP to enhance the AI model’s prediction accuracy, while also streamlining the mapping between the model’s input and output.
For the evaluation of AI-based cell-level mobility management, we should not only focus on the prediction accuracy of RSRP, such as the normalized mean squared error (NMSE) or RSRP difference between the actual RSRP and the predicted RSRP, but we also need to pay more attention to the following system-level KPIs which are important for HSR communications.
RLF rates: Radio link failure (RLF) occurs when the signal-to-interference-plus-noise ratio (SINR) stays consistently small for an extended duration. In the HSR scenarios, the ultra-high moving speed necessitates timely handovers to mitigate SINR attenuation and reduce RLF rates. The RLF rates can be used to evaluate the timeliness of the cell handover during train operation.
Handover failure (HOF) rates: Handover failure usually happens due to too late handover, too early handover and handover to a wrong cell. The HOF rates is calculated as the total number of HOFs divided by the sum of the total number of HOFs and the total number of successful handover. The HOF rates is one indicator on the reliability of the HSR communication system.
Data interruption: The time interval between the transmission of the last data packet from the source BS and the reception of the first data packet from the target BS is referred to as data interruption. The low data interruption guarantees the timeliness of signal quality for high-speed trains.
Table I summarizes the enhancement schemes for AI-based mobility management.
| Ref | Category | Network | Input | Output | Description |
| [9] | Beam-level | FC | The sub-6G channel vector | The optimal beam and the path state like LoS | The model utilizes the low-frequency sub-6G channel to learn the mmWave channel information. |
| [10] | FC | Location information, mobility vectors, SNR and current beam | The link connectivity within the next beam tracking phase | Using an AI model to predict whether there will be a connection loss before the next beam tracking procedure. | |
| [11] | CNN | LIDAR measurements, the coordinates of the BS and UE | The optimal beam and the path state (LoS or NLoS) | Generating a histogram based on the model’s input and feeding it into a CNN to learn the channel features. | |
| [7] | CNN+FC | RSRP of partial narrow beams or all wide beams | The direction of the transceiver beams | The model utilizes supervised and unsupervised learning based on CNN and FC. | |
| [8] | CNN+FC | The RSRP of all beams of each UE | The optimal beam index of each UE | The model adopts a hybrid beamforming architecture and CNN to achieve multi-user synchronous beam alignment. | |
| [2] | LSTM | The complete RSRP images | The optimal beam index | The multi-layer LSTM network is utilized for temporal beam prediction, followed by FC for feature extraction. | |
| [6] | LSTM | Channel state information (CSI) and the context vector | CSI and beam quality | The LSTM-based prediction model is embedded into the extended Kalman filter (EKF) algorithm. | |
| [5] | CNN+LSTM | The RSRP images of all beam pairs | The optimal beam index of transceiver | The CNN is utilized for feature extraction, followed by LSTM for temporal beam prediction. | |
| [4] | Conv-LSTM | The RSRP image of all wide beams | The RSRP image of all narrow beams | The model use sub-pixel convolution and convolutional LSTM realizing image reconstruction. | |
| Our | CS+AI | Channel information of compressed multi-beam | RSRP of all beams at the current instance | Learning the optimal measurement matrix and conv-LSTM network to reduce the cost of beam sweeping. | |
| [3] | Cell-level | CNN+LSTM | SNR of all beams and all cells | The same as the input | The model utilizes CNN to reduce the input dimension and LSTM network to predict beam quality of all cells. |
| [15] | CNN/LSTM | RSRP images of all beams of all cells | The best BS index, and path state (LoS or NLoS) | Using two AI models based on CNN and LSTM predicts the best BS index and path state in future instances. | |
| Our | CNN/LSTM | Compressed RSRP images | RSRP of all beams of all cells | This article uses two AI models to predict the RSRP in future instances based on compressed RSRP images. |
III-C AI Enhancement for HSR Communications
The traditional cell handover mechanism including 3 states is described in Fig. 4. In state 1, the UE measures the beam quality and sends the measurement report to the source BS. In this state, the UE remains connected to the source cell. When the mobility of the UE is low or the UE is moving within a macro cell, the above handover mechanism can sustain good performance. However, when the mobility of UE is high or the UE is moving across micro cells of high density, the traditional handover mechanism would result in unintended events, e.g., HOF, RLF, ping-pong phenomenon, throughput loss and too early/late handover. In HSR communications, the trains move at high speed, resulting in a sharp decline of the communication quality of the link between the UE and the source BS during state 1. Leveraging its robust environmental awareness and learning abilities, AI/ML technology can address the aforementioned issues.
One approach for AI-based cell-level mobility management is to exploit the beam-level outputs of different cells. In particular, AI models are deployed at the network side, and each cell employs a unique AI model to predict the future RSRP of this cell. The shortcomings of this cell-level mobility management approach lie in the high deployment and management overhead of AI models. Another approach is to use an AI model to predict the future RSRP of multiple adjacent cells of the same type based on RSRP measurement information. In this context, [15] utilized a multi-layer LSTM architecture to predict the optimal BS index as the candidate for handover. Moreover, [3] developed a CNN-based auto-encoder to reduce the data dimension, and then a LSTM network was used to conduct link quality prediction for each cell. In literature, the input into these AI models is the RSRP of beam pairs of all cells, leading to significant measurement overhead and computational complexity.
To this end, we consider two approaches to reduce the measurement overhead for cell-level mobility management. For the first approach, all beams from partially selected cells are used to predict the L3-RSRP for all cells. For the second approach, partially selected beams of each cell are used to predict the L3-RSRP for all cells. We utilize the partial L1-RSRP from the past six time instances to predict the L3-RSRP in the next four time instances. We employ CNN-based and LSTM-based models as shown in Fig. 5. The LSTM-based model comprises four LSTM layers and each layer has 168 hidden units. The features extracted by the LSTM network undergo further processing via an FC layer and a sigmoid function for feature mapping. To enhance the performance of the LSTM network, we adopt four identical networks to predict L3-RSRP for four future instances separately. For the CNN-based model, a set of regular 2D convolution layers with 12, 24 and 32 kernels are employed before feeding it into two FC layers. To avoid overfitting during the training process, max pooling layers are incorporated after each CNN layer.
The channel used in our experiments is generated by QuaDRiGa, where we consider the UMa scenario and center frequency 30 GHz. We deploy 7 BSs, and each BS is equipped with three half-wavelength spaced uniform plane arrays (UPAs) with sector orientation of 30, 150 and -90 degrees. We generate four datasets corresponding to various UE speeds of 60 km/h, 120 km/h, 350 km/h, and 500 km/h. Each dataset comprises 50,000 samples, which are split into training, validation, and test sets in a 3:1:1 ratio. For the proposed two approaches with equidistant sampling, L1-RSRP of 50% beams in each cell or all beams in 50% cells are selected as input to the AI model. The batch size is 256 and the maximum epoch is 800. The learning rate dynamically changes from to , utilizing the adaptive moment estimation optimizer. The RSRP difference is one evaluation metric of prediction accuracy for cell-level mobility management, which reflects the disparity between the predicted L3-RSRP and the actual L3-RSRP of all cells. In addition, we also consider the RLF rates as the system-level evaluation metric to illustrate the overall handover performance of the communication system.
The RSRP difference between the AI predictions and the actual measurements is shown in Fig. 6. The experimental results show that both CNN-based and LSTM-based models show excellent cell-level temporal prediction performance. Table II shows the significantly reduced RLF rates in AI-assisted cell handover processes. Remarkably, the proposed two approaches to reduce the measurement overhead for cell-level beam management, labeled as Part_Cell and Part_Beam, achieve comparable RLF performance with the traditional approach using all beam measurements of all cells, labeled as All_Beam_Cell, while we can save 50% beam measurement overhead for HSR communications. Furthermore, as shown in Table II, the flops and total parameters of the CNN-based model are much lower than those of the LSTM-based model.
| UE Speed | Flops | Total Params | ||||
| 60km/h | 120km/h | 350km/h | 500km/h | |||
| Non-AI | 24.61% | 27.75% | 39.06% | 39.91% | ||
| CNN_ALL_Beam_Cell | 5.08% | 9.44% | 10.07% | 10.07% | 33.85M | 28.66M |
| CNN_Part_Cell | 5.39% | 9.39% | 9.84% | 10.33% | 5.59M | 3.61M |
| CNN_Part_Beam | 5.36% | 9.26% | 10.23% | 10.45% | 4.82M | 2.51M |
| LSTM_ALL_Beam_Cell | 5.08% | 9.57% | 10.06% | 10.22% | 391.87M | 66.02M |
| LSTM_Part_Cell | 5.32% | 9.59% | 10.07% | 10.43% | 354.46M | 58.46M |
| LSTM_Part_Beam | 5.08% | 9.58% | 9.87% | 10.22% | 348.52M | 58.80M |
IV Opportunities and Challenges of AI-Based Mobility Management for HSR Communications
IV-A Dynamic Interference Management
Due to the high mobility in HSR communications, the randomness of inter-cell interference constitutes a crucial challenge. Specifically, mmWave beams from neighboring cells directed at local terminals can cause severe inter-cell interference, while frequent cell handovers and beam switching exacerbate the randomness of this interference. Therefore, agile interference management is crucial to ensure reliable service in HSR communications. As a potential solution, graph neural networks (GNNs) can intrinsically match the topology of wireless networks and can be used to effectively handle dynamic interference management.
IV-B Multi-Criteria Joint Decision-Making Mechanism for Mobility Management
In HSR environments, the rapid movement necessitates more adaptable mobility management strategies to effectively handle frequent environmental changes. Relying solely on RSRP variations is insufficient for ensuring communication stability during handover decisions. Therefore, a multi-criteria decision-making mechanism is imperative. This mechanism could integrate factors, such as signal interference levels, network load conditions, terminal velocity, and direction of movement, along with RSRP variations as inputs for AI models to obtain better handover decisions. This would greatly improve the accuracy and reliability, ensuring seamless communication for high-speed trains.
IV-C Model Deployment and Compression
In practical scenarios, the AI-assisted mobility management methods require high energy dissipation due to the high computing complexity, so it is more meaningful to deploy these AI models on the network side. In this case, the UE reports measurement data to the source BS, which feeds the measurements into the AI/ML models for mobility management. Moreover, the transmission entails extra latency while the unavoidable transmission errors, occurring in reporting these measurements, affect the performance of AI models. Then, it would be beneficial to deploy AI models on the UE side. We can consider using CS or training lightweight AI models to reduce the computational complexity and hardware requirements for UEs.
IV-D Model Generalization Design
AI/ML models designed for specific scenarios would usually have poor generalization performance. Altering the settings in communication scenarios, e.g., BS height, channel model, cell size and UE speed, might lead to a significant deterioration of the model performance. Therefore, one challenge for AI-based mobility management for HSR communications is the design of AI models with good scenario generalization properties. Generally, there exists a trade-off between model generalization and model performance. Improving the generalization of the model, e.g., training the model with a mixed datasets from multiple scenarios, will introduce difficulties for the training of AI/ML models, as more complex and diverse knowledge are embedded in the data. As a result, the balance between model generalization and prediction accuracy according to the actual application scenario needs to be thoroughly examined in future discussions.
V Conclusion
In this article, we investigated the fundamental issue of AI-based beam-level and cell-level mobility management for HSR communications, including the use cases, inputs, outputs and KPIs of AI models. For beam-level mobility management in HSR communications, a hybrid approach, combining AI and CS, demonstrated superior performance compared to conventional AI-based methods with the same measurement overhead. For cell-level mobility management in HSR communications, we unveiled the significant gain of AI-based methods for cell handover, while the measurement overhead can be reduced by using partial beam measurements, without compromise of the RLF rates. At last, the opportunities and challenges on AI-based mobility management are briefly discussed for further studies.

### v2

@@
-AI-Based Beam-Level and Cell-Level Mobility Management for High Speed Railway Communications
+AI-Driven Mobility Management for High-Speed Railway Communications: Compressed Measurements and Proactive Handover
 Abstract
@@
 The high speed of trains creates rapidly time-varying wireless channels, increases the signaling overhead, and reduces the system throughput, making it difficult to meet the growing and stringent needs of HSR applications.
-In this article, we explore artificial intelligence (AI)-based beam-level and cell-level mobility management suitable for HSR communications, including the use cases, inputs, outputs, and key performance indicators (KPI)s of AI models.
-Particularly, in comparison to traditional down-sampling spatial beam measurements, we show that the compressed spatial multi-beam measurements via compressive sensing lead to improved spatial-temporal beam prediction.
-Moreover, we demonstrate the performance gains of AI-assisted cell handover over traditional mobile handover mechanisms.
-In addition, we observe that the proposed approaches to reduce the measurement overhead achieve comparable radio link failure performance with the traditional approach that requires all the beam measurements of all cells, while the former methods can save 50% beam measurement overhead.
+In this article, we explore artificial intelligence (AI)-based beam-level and cell-level mobility management suitable for HSR communications.
+Particularly, we propose a compressed spatial multi-beam measurements scheme via compressive sensing for beam-level mobility management in HSR communications.
+In comparison to traditional down-sampling spatial beam measurements, this method leads to improved spatial-temporal beam prediction accuracy with the same measurement overhead.
+Moreover, we propose a novel AI-based proactive handover scheme to predict handover events and reduce radio link failure (RLF) rates in HSR communications.
+Compared with the traditional event A3-based handover mechanism, the proposed approach significantly reduces the RLF rates which saves 50% beam measurement overhead.
 Index Terms:
-Artificial intelligence, beam-level mobility management, cell-level mobility management.I Introduction
-High speed railway (HSR) provides people with safe, fast, comfortable and economical modes of transportation, and has been widely developed around the world.
+Artificial intelligence, compressive sensing, mobility management.I Introduction
+High-speed railway (HSR) provides people with safe, fast, comfortable and economical modes of transportation, and has been widely developed around the world.
 Nevertheless, mobile communication systems are pivotal for ensuring rail safety, operations, maintenance, and delivering passenger information services.
-Emerging applications, such as railway multimedia dispatching, high-definition video monitoring on board and trackside, have posed stricter requirements on the reliability, data rate, and spectrum efficiency of the next-generation, dedicated mobile communication systems for HSR.
+Emerging applications, such as railway multimedia dispatching and high-definition video monitoring on board and trackside, have posed stricter requirements on the reliability, data rate, and spectral efficiency of the next-generation, dedicated mobile communication systems for HSR.
 However, the high speed of trains entails formidable challenges due to the time-varying nature of wireless channels in HSR communications.
@@
 Thanks to the abundant spectrum available at higher frequencies, communication systems using the millimeter wave (mmWave) technology avail of faster data transmission speeds and reduced latency compared to traditional low-frequency communications.
-For HSR communications, mmWave technology underpins increased data transmission rates that meet the demands of onboard applications and services, and provides low latency that is indispensable for real-time communication requirements.
+For HSR communications, mmWave frequency bands can underpin increased data transmission rates that meet the demands of onboard applications and services, and provides low latency that is indispensable for real-time communication requirements.
 However, mmWave communication systems suffer from significant path loss, which restricts the effective range of transmission.
@@
 However, the high mobility of trains entails pressing challenges in mobility management for HSR communications.
-While being connected, a train has to address two types of mobility, i.e., the cell-level and beam-level mobility as specified in the 3rd Generation Partnership Project (3GPP) Release 17 for 5G systems[1].
+While being connected, a train has to address two types of mobility, i.e., the beam-level and cell-level mobility as specified in the 3rd Generation Partnership Project (3GPP) Release 17 for fifth generation (5G) systems[1].
 Beam-level mobility involves dynamically switching beams in which the train establishes a link path to a new beam within the same cell, ensuring communication with high performance.
@@
 As trains move from one coverage zone to another, it is crucial to have a smooth cell handover to sustain continuous connectivity experience.
-The management of mobility at the beam and cell level in 5G is based on the use of periodic channel state information reference signals (CSI-RS) and synchronization signal blocks (SSBs).
-Unfortunately, the resulting large latency and high overhead are prohibitive for HSR communications, and, hence, innovative approaches are urgently needed to address these limitations.
-Deep neural networks (DNNs) have found extensive applications in areas, such as computer vision and natural language processing, owing to their exceptional ability to extract features from intricate and high-dimensional data.
-In recent years, DNNs have also proven their effectiveness in capturing the complex and nonlinear characteristics of wireless channels to facilitate beam-level mobility management.
-For example, a multi-layer long short term memory (LSTM) network was utilized to predict the optimal beam change for future instances [2].
-Studies on advanced artificial intelligence (AI)-based cell-level mobility management are relatively few.
-For example, [3] utilized convolutional neural network (CNN)-based auto-encoders for dimension reduction and LSTM networks to predict the appropriate base station (BS) for cell handover.
-Recently, the 3GPP 5G NR organization has launched studies on AI-based cell-level mobility management for highly reliable handovers in Release 19.
-However, these AI-based mobility management methods fail to exploit the specific characteristics in HSR communications.
-For instance, trains typically maintain stable high-speed motion on fixed tracks and have different speed levels to meet different needs.
-Mobility management methods for HSR communications are expected to leverage case-specific characteristics in the design phase or, alternatively, learn from data to yield improved performance in a data-driven manner.
-In this article, we present the latest progress in AI-based beam-level and cell-level mobility management suitable for HSR communications.
+Beam-level mobility management in 5G relies on periodic channel state information reference signals and synchronization signal blocks for beam sweeping, measurement, and reporting.
+However, frequent beam sweeping introduces significant latency and high overhead[2], making it impractical for HSR communications.
+Cell-level mobility management primarily relies on an event A3-based handover process, encompassing handover trigger, preparation, and execution.
+Before handover completion, the UE continues to communicate with the source base station (BS), often resulting in a sharp decline in the communication quality.
+Additionally, with a fixed handover period, the link quality between the high-speed terminal and source BS deteriorates faster, increasing the risk of handover failure.
+Hence, innovative handover mechanisms are urgently needed to address these limitations.
+In recent years, deep neural networks (DNNs) have proven their effectiveness in capturing the complex and nonlinear characteristics of wireless channels to enhance mobility management.
+For example, in [2] the optimal beam for future instances was predicted by using a multi-layer long short term memory (LSTM) network.
+The appropriate BS for cell handover was predicted by using convolutional neural network (CNN)-based auto-encoders and LSTM networks in [3].
+However, these methods either measure all beams, resulting in high overhead, or measure a part of all beams, resulting in notable performance loss.
+Additionally, existing AI-based cell-level mobility management methods focus solely on the prediction accuracy of the reference signal received power (RSRP), ignoring the communication quality degradation resulting from the reactive nature of the event A3-based handover mechanism.
+Recently, the 3GPP 5G-Advanced New Radio (NR) standardization has launched studies on AI-based mobility management for highly reliable handovers in Release 19.111https://www.3gpp.org/ftp/tsg_ran/TSG_RAN/TSGR_102/Docs/RP-234055.zip Mobility management methods for HSR communications are expected to leverage case-specific characteristics in the design phase or, alternatively, learn from data to yield improved performance in a data-driven manner.
+In this article, we provide two AI-based beam-level and cell-level mobility management approaches for HSR communications.
 In Section II, we reveal the potential of AI-based spatial-temporal beam prediction for HSR communications.
-Particularly, in comparison to the traditional down-sampling spatial beam measurements, we show that the compressed spatial multi-beam measurements via compressive sensing (CS) can lead to improved spatial-temporal beam prediction in HSR communications.
-In Section III, we investigate AI-based cell-level mobility management for HSR scenarios, including use cases, the input/output of the AI model, evaluation indicators, and enhanced schemes.
-Particularly, we demonstrate the performance gains of AI-assisted cell handover over traditional cell handover mechanisms for HSR scenarios.
-Section V introduces the challenges and opportunities encountered in AI-based mobility management, followed by our conclusions in Section VI.
-II AI-Based Beam-Level Mobility Management for HSR Communications
-II-A Use Cases in HSR Communications
-Beam-level mobility enables a user equipment (UE) to communicate with a BS via the optimal beam pair, which involves low-level beam operations rather than handover.
+Particularly, by leveraging the sparsity and time-domain correlation of channels in HSR communications, we propose compressed spatial multi-beam measurements via compressive sensing (CS) to improve the spatial-temporal beam prediction accuracy in HSR communications.
+In Section III, we investigate AI-based cell-level mobility management for HSR scenarios and propose a novel AI-assisted cell handover mechanism, performing handover proactively before the signal quality deteriorates by AI-based handover event prediction.
+Section V introduces the challenges and opportunities encountered in AI-based mobility management for HSR communications, followed by our conclusions in Section VI.
+A summary of our overview of AI-based mobility enhancement schemes is given in Table I.
+II AI-Based Compressed Multi-Beam Measurement for HSR Communications
+II-A AI-Based Beam-Level Mobility Management in HSR Communications
+When a user equipment (UE) performs a beam switch within the same cell, it initiates beam-level mobility.
+This process, which includes beam sweeping, beam measurement, beam reporting, and beam determination is managed within the cell, where the physical (PHY) and media access control (MAC) layers are sufficient to handle it, thereby avoiding the need for radio resource control (RRC) operations.
+The detailed beam-level mobility management process is shown in Fig.
+2.
+Beam-level mobility management enhancement involves three types of beam search methods: traditional methods including exhaustive search and hierarchical search, advanced numerical methods, and AI-based methods.
+The exhaustive beam search method examines all the candidate beams in the predefined codebook one by one to obtain beam quality information.
+The hierarchical beam search method provides a more efficient alternative, which first searches within a wide beam range and then further searches for a narrow beam within the angular range of the selected wide beam.
+However, traditional beam search algorithms suffer from a large measurement overhead, which is not suitable for HSR communications which require frequent beam switching.
+In addition, enhancement techniques, based on numerical methods have been proposed.
+More specifically, [4] exploited the tree search method for efficient beam training based on the hierarchical codebook, whilst [5] proposed an optimal beam training model based on Markov’s decision process.
+These numerical approaches mainly target general communication environments, making it challenging to tailor them for particular communication scenarios.
 In HSR communications, with some coarse-grained beam quality measurements of the serving cell, AI can be utilized to estimate the finely-grained optimal beam pair in the current instance or predict the optimal beam pair in future instances, which leads to reduced measurement overhead and enhanced throughput.
@@
 Spatial beam prediction exploits spatial correlations among adjacent beams to estimate the quality of all beams using incomplete beam data.
-CNNs and networks with fully connected (FC) layers have been commonly employed for this purpose [7, 8, 9].
-Due to the high mobility in HSR communications, the extensive signaling overhead and the time required for frequent beam sweeping become prohibitive.
-For these reasons, temporal beam prediction offers greater benefits in terms of resource conservation and stable performance in HSR communications.
-II-B Inputs and Outputs in HSR Communications
-Different tasks in beam-level mobility management require distinct input/output into/of the AI models in HSR communications.
-According to the ongoing discussions within 3GPP, there are two vital concepts for beam management, namely Set A and Set B.
-Set A comprises all narrow beams from the predefined codebook, while Set B encompasses the beams that have been measured and are utilized as inputs into the AI models.
-Various alternative methods exist with respect to the relationship between Set A and Set B.
-For instance, the AI model could receive the input of reference signal received power (RSRP) measurements of partial narrow beams, making Set B as a subset of Set A.
+CNNs and networks with fully connected (FC) layers have been commonly employed for this purpose [9, 10, 11, 12].
+Due to the high mobility in HSR communications, the signaling overhead and the time required for frequent beam sweeping become prohibitive.
+For these reasons, temporal or spatial-temporal beam prediction offers greater benefits in terms of resource conservation and stable performance of HSR communications.
+Different tasks in beam-level mobility management require distinct inputs to the AI models in HSR communications.
+In order to instigate further discussion, there are two vital concepts for beam-level mobility management, namely Set A and Set B.222https://www.3gpp.org/ftp/tsg_ran/WG1_RL1/TSGR1_109-e/Inbox/drafts/9.2.3.2/R1-220xxxx%20%5B109-e-R18-AIML-06%5D-Summary-1-v022-MediaTek_Intel.docx Set A comprises all narrow beams from the predefined codebook, while Set B encompasses the beams that have been measured and are utilized as inputs into the AI models.
+Various alternative model input methods exist with respect to the relationship between Set A and Set B.
+For example, the AI model could receive the input of RSRP measurements of partial narrow beams, making Set B a subset of Set A.
 This is the most commonly used method in HSR communications, which greatly reduces the measurement overhead.
 Another option is to utilize RSRP measurements of wide beams to estimate the quality of narrow beams, where Set B and Set A are distinct.
-In addition, Set B can be the same as Set A, whenever the RSRP measurements of all narrow beams are the inputs into the AI models.
-Note that the first two types of inputs are employed in temporal and spatial beam prediction models, while the last is only for temporal beam prediction.
-Instead of using RSRP, the complex received signal can be used as input into the AI model, which will be discussed in the CS-based method in the next subsection.
-In addition, supplementary auxiliary information can be incorporated to improve the prediction accuracy for HSR communications.
-In [10], the position of the UE, its mobility vector, the signal-to-noise ratio (SNR), and the beam index currently in use were utilized to forecast the state of the link connection.
-Furthermore, measurements from light detection and range (LIDAR) and various sensors, images of the environment taken by cameras, and information from low-frequency channels could also be employed as extra inputs into the AI model to enhance the prediction accuracy.
-There are two types of outputs in the AI model design for beam-level mobility management.
-The most common outputs are weights associated with different beams, where a higher weight suggests a better beam.
-Additionally, line-of-sight (LoS) is the primary mode of signal propagation in HSR communications and beam-selection is easier in the LoS setting.
-Therefore, the AI model can be utilized to output the estimated state of each path, i.e., LoS or non-LoS, to enhance the beam prediction [11, 9].
-II-C AI Enhancement for HSR Communications
-Beam-level mobility management enhancement involves three types of beam search methods: traditional methods including exhaustive search and hierarchical search, advanced numerical methods, and AI-based methods.
-The exhaustive beam search method examines all the candidate beams in the predefined codebook one by one to obtain beam quality information.
-The hierarchical beam search method provides a more efficient alternative, which first searches within a wide beam range and then further searches for a narrow beam within the angular range of the selected wide beam.
-However, traditional beam search algorithms suffer from a large training overhead, which is not suitable for HSR communications which require frequent beam switching.
-In addition, enhancement techniques, based on numerical methods have been proposed.
-More specifically, [12] exploited the tree search method for efficient beam training based on the hierarchical codebook, whilst [13] proposed an optimal beam training model based on Markov’s decision process.
-These numerical approaches mainly target general communication environments, making it challenging to tailor them for particular communication scenarios.
+In addition, whenever the RSRP measurements of all narrow beams are the inputs into the AI models, Set B can be the same as Set A.
+The first two types of inputs, where the inputs to the model are RSRP measurements of some narrow beams and all wide beams, are designed for temporal and spatial beam prediction models, aiming to reduce the measurement overhead by lowering the frequency of beam measurements, which makes them more suitable for latency-sensitive applications.
+In contrast, the last type of input relies on measuring the RSRP of all candidate beams to conduct temporal beam prediction, achieving the higher model prediction accuracy.
+This method is well-suited for applications that demand exceptionally high accuracy in prediction models and are less sensitive to latency.
+Instead of using RSRP, supplementary auxiliary information, like the position of the UE, its mobility vector, the signal-to-noise ratio (SNR), and the beam index can currently be utilized to provide broader spatial context to forecast the state of the link connection [13].
+II-B AI-Based Compressed Multi-Beam Measurement Scheme
 In HSR communications, every train follows a predetermined path and performs the same routes periodically.
@@
 The compressed channel information is fed into the remaining neural network to reconstruct the complete RSRP information of all candidate beams for a future instance.
+In the remaining module, a sub-pixel convolution layer and a 2D upsampling layer are applied before data is fed into the convolutional LSTM (Conv-LSTM) layers.
+These two layers learn to upsample low-resolution features from limited beam measurements, mapping them to high-resolution feature maps of the complete beam measurements.
+Then, the initial estimates of the complete beam measurement map are fed into the Conv-LSTM layers to calculate the temporal variations by leveraging historical data and local correlations.
+Finally, the network outputs a complete RSRP map for an upcoming time slot.
 As a learnable parameter, the CS measurement matrix is jointly optimized during the training of the model.
-In the inference phase, the trained measurement matrices are applied at the transceivers and the compressed measurements are fed into the trained neural network to conduct beam prediction.
-Compared with the traditional sweeping method, the multi-beam compression offers to the beam prediction network richer information with the same number of beam measurements.
 The structure of the proposed AI model for beam-level mobility management, namely CSAI, is illustrated in Fig.
-2(d).
-Now, we compare the performance of AI-based spatial-temporal beam prediction methods, whose structures are shown in in Fig.
-2.
+3(d).
+In the inference phase, the trained measurement matrices are applied at the transceivers while the compressed channel measurements are fed into the trained neural network to conduct beam prediction.
+Meanwhile, the model can undergo periodic retraining or continuous updates through online learning.
+Detailed deployment information is shown in Fig.
+2 Compared with the traditional sweeping method, the multi-beam compression offers to the beam prediction network richer information with the same number of beam measurements.
+Now, we compare the performance of AI-based spatial-temporal beam prediction methods, whose structures are shown in Fig.
+3.
 The AI-based methods in Fig.
-2(a) [6] and Fig.
-2(b) [7] employ the traditional LSTM and CNN architecture, respectively, to predict the high-resolution beams.
+3(a) [8] and Fig.
+3(b) [9] employ the traditional LSTM and CNN architecture, respectively, to predict the high-resolution beams.
 Compared with the CSAI in Fig.
-2(d), the AI-based method in Fig.
-2(c) [4] has the same network structure except for the beam measurement module.
-In the experiments, we consider the UMa scenario for outdoor 30 GHz mmWave communications, while the UE follows a straight-line trajectory which aligns with the HSR scenario.
+3(d), the AI-based method in Fig.
+3(c) [6] has the same network structure except for the beam measurement module.
+In the experiments, the channel used in our experiments is generated by QuaDRiGa, and we consider the UMa scenario for outdoor 30 GHz mmWave communications, while the UE follows a straight-line trajectory which aligns with the HSR scenario.
 Four distinct datasets, representing UE speeds of 60km/h, 120km/h, 350km/h and 500km/h, are utilized in the evaluation.
@@
 The learning rate changes dynamically from to , utilizing the adaptive moment estimation optimizer.
-We consider the prediction accuracy for the best beam with a compression ratio111The compression ratio is defined as the ratio of the number of input beam measurements to the number of all beam measurements.
+We consider the prediction accuracy for the best beam with a compression ratio333The compression ratio is defined as the ratio of the number of input beam measurements to the number of all beam measurements.
 of .
-The simple non-AI approach straightforwardly uses the best one of the partially measured present beams as the prediction.
+The simple non-AI approach straightforwardly uses the best one out of the partially measured present beams as the prediction.
+For three traditional AI-based beam prediction methods shown in Fig.
+3(a) (b) (c), the model’s input uses equal interval sampling from predefined codebook to obtain beam measurement information with a compression ratio of .
 As shown in Fig.
-3, the AI-based spatial-temporal beam prediction methods outperform significantly the traditional non-AI method.
-As the speed of the movement increases, all methods generally show an anticipated decrease in the prediction accuracy.
+4, the AI-based spatial-temporal beam prediction methods outperform significantly the traditional non-AI method.
+As the speed of the movement increases, all methods generally show an anticipated degradation in the prediction accuracy.
 The CSAI demonstrates superior performance for all the cases.
 At the speed of 500km/h, the CSAI is still able to yield a prediction accuracy of 83.76%.
-III AI-Based Cell-level Mobility Management in HSR Communications
-III-A Use Cases in HSR Communications
+The Conv-LSTM network shown in Fig.
+3(c) requires 3.84M flops and 0.53M parameters, whereas the proposed CSAI model utilizes 3.88M flops and 0.136M parameters.
+This demonstrates that the proposed model improves prediction performance with a negligible increase in computational cost compared to the fixed down-sampling beam pattern approach.
+III AI-Based Proactive Handover Mechanism for HSR Communications
+III-A AI-Based Cell-Level Mobility Management in HSR Communications
 Cell-level mobility management deals primarily with cell handovers, initiating the transition of a user’s session to a new cell when the signal strength from the serving cell falls below a predefined threshold.
-Given the rapid movement in HSR communications, cell handovers occur frequently, and failures in these handovers would unavoidable compromise the user experience.
-As specified in 3GPP, the cell handover procedure relies on concrete criteria, such as event A3, which will be triggered when the difference between the layer 3 RSRP (L3-RSRP)222L3-RSRP represents the cell-level RSRP, obtained by averaging or maximizing the RSRP of all beam pairs in each cell.
+Given the rapid movement in HSR communications, cell handovers occur frequently, and failures in these handovers would unavoidably compromise the user experience.
+Handover decision is performed by the source cell based on UE measurement.
+UE measurement should be met by certain criteria, for example, event A3 which will be triggered when the difference between the layer 3 RSRP (L3-RSRP)444L3-RSRP represents the cell-level RSRP, obtained by averaging or maximizing the RSRP of all beam pairs in each cell.
 measurement of the target cell and the L3-RSRP measurement of the source cell exceeds a predefined offset.
 If the above criteria are met during the time to trigger (TTT) duration, the UE reports the measurements and the source cell initiates the handover preparation phase.
-Afterwards, the cell handover procedure initiates the execution after the UE receives the handover command message, whose task is to provide radio resource control (RRC) reconfiguration message to the target cell.
+Afterwards, the cell handover procedure initiates the execution after the UE receives the handover command message, whose task is to provide RRC reconfiguration message to the target cell.
 The specific cell handover process is shown in Fig.
-4.
-During state 1, the UE must sweep across all beams to acquire the L3-RSRP of adjacent cells for the TTT duration, leading to signal quality deterioration in the service cell, which are even more pronounced in HSR scenarios.
-To maintain a stable communication quality for HSR communications, the most common use case for AI-based cell-level mobility management is to predict the RSRP of all adjacent cells in future TTT instances.
-The AI-based cell handover process is illustrated in Fig.
-4, where option 1 and option 2 represent different deployment methods of AI models.
-In this case, one option for the source cell is to immediately execute a cell handover based on the AI model’s predicted target cell.
-Alternatively, the source cell can monitor the signal quality of both the target and source cells over a period of time and initiate the handover if some handover conditions are still met.
-This approach can effectively reduce the ping-pong handovers and enhance the communication stability, while it requires more handover time for signal detection.
-In HSR scenarios, where trains move in straight lines and the probability of ping-pong handovers is low, the delay introduced by extended handover time, resulting in significant communication quality degradation, becomes intolerable.
-Therefore, the first approach is better suited for cell-level mobility management in HSR communications.
-III-B Inputs, Outputs, and KPIs in HSR Communications
-For the HSR scenario, we explore various model inputs and outputs tailored to its specific requirements.
-One input option for AI models is the L3-RSRP of each cell at historical instances, with the output being either the L3-RSRP of each cell or the L3-RSRP of the top cells at future TTT instances.
-The dimension of L3-RSRP information is lower than that of L1-RSRP.
-However, acquiring L3-RSRP, which requires averaging data over multiple time slots, introduces extra latency which is not desired in HSR communications.
-Another viable approach involves using layer 1 RSRP (L1-RSRP)333L1-RSRP represents the beam-level RSRP, which is the RSRP of each beam pair of each cell.
-as both the input and the output of the AI models.
-The AI model output is then filtered to derive the L3-RSRP, which is used to execute cell handovers.
-For this option, the output L1-RSRP of AI model can be applied to both beam-level and cell-level mobility management.
-Nevertheless, L1-RSRP encompasses extensive information that exceeds the requirements for handover, while the correlation between the input and output L1-RSRP is intricate, thereby amplifying the learning challenge for the AI models.
-Moreover, leveraging L1-RSRP to forecast L3-RSRP via AI can be an advantageous strategy.
-This method reaps the benefits of the above two approaches, making use of the abundant data in L1-RSRP to enhance the AI model’s prediction accuracy, while also streamlining the mapping between the model’s input and output.
-For the evaluation of AI-based cell-level mobility management, we should not only focus on the prediction accuracy of RSRP, such as the normalized mean squared error (NMSE) or RSRP difference between the actual RSRP and the predicted RSRP, but we also need to pay more attention to the following system-level KPIs which are important for HSR communications.
-RLF rates: Radio link failure (RLF) occurs when the signal-to-interference-plus-noise ratio (SINR) stays consistently small for an extended duration.
-In the HSR scenarios, the ultra-high moving speed necessitates timely handovers to mitigate SINR attenuation and reduce RLF rates.
-The RLF rates can be used to evaluate the timeliness of the cell handover during train operation.
-Handover failure (HOF) rates: Handover failure usually happens due to too late handover, too early handover and handover to a wrong cell.
-The HOF rates is calculated as the total number of HOFs divided by the sum of the total number of HOFs and the total number of successful handover.
-The HOF rates is one indicator on the reliability of the HSR communication system.
-Data interruption: The time interval between the transmission of the last data packet from the source BS and the reception of the first data packet from the target BS is referred to as data interruption.
-The low data interruption guarantees the timeliness of signal quality for high-speed trains.
-Table I summarizes the enhancement schemes for AI-based mobility management.
-| Ref | Category | Network | Input | Output | Description |
-| [9] | Beam-level | FC | The sub-6G channel vector | The optimal beam and the path state like LoS | The model utilizes the low-frequency sub-6G channel to learn the mmWave channel information. |
-| [10] | FC | Location information, mobility vectors, SNR and current beam | The link connectivity within the next beam tracking phase | Using an AI model to predict whether there will be a connection loss before the next beam tracking procedure.
-| |
-| [11] | CNN | LIDAR measurements, the coordinates of the BS and UE | The optimal beam and the path state (LoS or NLoS) | Generating a histogram based on the model’s input and feeding it into a CNN to learn the channel features.
-| |
-| [7] | CNN+FC | RSRP of partial narrow beams or all wide beams | The direction of the transceiver beams | The model utilizes supervised and unsupervised learning based on CNN and FC. | |
-| [8] | CNN+FC | The RSRP of all beams of each UE | The optimal beam index of each UE | The model adopts a hybrid beamforming architecture and CNN to achieve multi-user synchronous beam alignment. | |
-| [2] | LSTM | The complete RSRP images | The optimal beam index | The multi-layer LSTM network is utilized for temporal beam prediction, followed by FC for feature extraction. | |
-| [6] | LSTM | Channel state information (CSI) and the context vector | CSI and beam quality | The LSTM-based prediction model is embedded into the extended Kalman filter (EKF) algorithm. | |
-| [5] | CNN+LSTM | The RSRP images of all beam pairs | The optimal beam index of transceiver | The CNN is utilized for feature extraction, followed by LSTM for temporal beam prediction. | |
-| [4] | Conv-LSTM | The RSRP image of all wide beams | The RSRP image of all narrow beams | The model use sub-pixel convolution and convolutional LSTM realizing image reconstruction. | |
-| Our | CS+AI | Channel information of compressed multi-beam | RSRP of all beams at the current instance | Learning the optimal measurement matrix and conv-LSTM network to reduce the cost of beam sweeping.
-| |
-| [3] | Cell-level | CNN+LSTM | SNR of all beams and all cells | The same as the input | The model utilizes CNN to reduce the input dimension and LSTM network to predict beam quality of all cells. |
-| [15] | CNN/LSTM | RSRP images of all beams of all cells | The best BS index, and path state (LoS or NLoS) | Using two AI models based on CNN and LSTM predicts the best BS index and path state in future instances.
-| |
-| Our | CNN/LSTM | Compressed RSRP images | RSRP of all beams of all cells | This article uses two AI models to predict the RSRP in future instances based on compressed RSRP images. |
-III-C AI Enhancement for HSR Communications
-The traditional cell handover mechanism including 3 states is described in Fig.
-4.
+5.
 In state 1, the UE measures the beam quality and sends the measurement report to the source BS.
 In this state, the UE remains connected to the source cell.
-When the mobility of the UE is low or the UE is moving within a macro cell, the above handover mechanism can sustain good performance.
-However, when the mobility of UE is high or the UE is moving across micro cells of high density, the traditional handover mechanism would result in unintended events, e.g., HOF, RLF, ping-pong phenomenon, throughput loss and too early/late handover.
+When the mobility of the UE is low or the UE is moving within a macro cell, the traditional handover mechanism can sustain good performance.
+However, when the mobility of UE is high or the UE is moving across micro cells of high density, the traditional handover mechanism would result in unintended events, e.g., handover failure, radio link failure (RLF), ping-pong phenomenon, throughput loss.
 In HSR communications, the trains move at high speed, resulting in a sharp decline of the communication quality of the link between the UE and the source BS during state 1.
@@
 The shortcomings of this cell-level mobility management approach lie in the high deployment and management overhead of AI models.
-Another approach is to use an AI model to predict the future RSRP of multiple adjacent cells of the same type based on RSRP measurement information.
+Another approach which is more suitable for HSR scenarios is to use an AI model to predict the future RSRP of multiple adjacent cells or the future optimal BS index based on historical RSRP measurement information.
 In this context, [15] utilized a multi-layer LSTM architecture to predict the optimal BS index as the candidate for handover.
 Moreover, [3] developed a CNN-based auto-encoder to reduce the data dimension, and then a LSTM network was used to conduct link quality prediction for each cell.
-In literature, the input into these AI models is the RSRP of beam pairs of all cells, leading to significant measurement overhead and computational complexity.
-To this end, we consider two approaches to reduce the measurement overhead for cell-level mobility management.
-For the first approach, all beams from partially selected cells are used to predict the L3-RSRP for all cells.
-For the second approach, partially selected beams of each cell are used to predict the L3-RSRP for all cells.
+For the HSR scenario, we also explore various model inputs and outputs tailored to its specific requirements.
+One input option for AI models is the L3-RSRP of each cell, with the output being either the L3-RSRP of each cell or the L3-RSRP of the top cells.
+The dimension of L3-RSRP information is lower than that of layer 1 RSRP (L1-RSRP)555L1-RSRP represents the beam-level RSRP, which is the RSRP of each beam pair of each cell.
+However, acquiring L3-RSRP, which requires averaging data over multiple time slots, introduces extra latency which is not desired in HSR communications.
+Another viable approach involves using L1-RSRP as both the input to and the output of the AI models.
+The AI model output is then filtered to derive the L3-RSRP, which is used to execute cell handovers.
+For this option, the output L1-RSRP of the AI model can be applied to both beam-level and cell-level mobility management.
+Nevertheless, L1-RSRP encompasses extensive information that exceeds the requirements for handover, while the correlation between the input and output L1-RSRP is intricate, thereby amplifying the learning challenge for the AI models.
+Moreover, leveraging L1-RSRP to forecast L3-RSRP via AI can be an advantageous strategy.
+This method reaps the benefits of the above two approaches, making use of the abundant data in RSRP to enhance the AI model’s prediction accuracy, while also streamlining the mapping between the model’s input and output.
+Ref | Category | Network | Input | Output | Description | KPIs |
+[11] | Beam-level | FC | The sub-6G channel vector | The optimal beam and the path state like LoS | Utilizing the low-frequency sub-6G channel to predict. | Top-1 accuracy, spectral efficiency |
+[13] | FC | Location information, mobility vectors, SNR and current beam | The link connectivity within the next beam tracking phase | Predicting whether there will be a connection loss before the next beam tracking procedure.
+| Throughput | |
+[12] | CNN | LIDAR measurements, the coordinates of the BS and UE | The optimal beam and the path state: line-of-sight (LoS) or non line-of-sight (NLoS) | Generating a histogram based on the model’s input to learn the channel features.
+| Accuracy for LoS detection | |
+[9] | CNN+FC | RSRP of partial narrow beams or all wide beams | The direction of the transceiver beams | Utilizing supervised and unsupervised learning based on CNN and FC. | Achievable rate | |
+[10] | CNN+FC | The RSRP of all beams of each UE | The optimal beam index of each UE | Adopting a CNN to achieve multi-user synchronous alignment. | Spectral efficiency | |
+[2] | LSTM | The complete RSRP images | The optimal beam index | The multi-layer LSTM network and FC are utilized. | RSRP difference,Top-K accuracy | |
+[8] | LSTM | Channel state information (CSI) and the context vector | CSI and beam quality | The LSTM network and extended Kalman filter (EKF) are used. | Achievable rate | |
+[7] | CNN+LSTM | The RSRP images of all beam pairs | The optimal beam index of transceiver | Utilizing CNN for feature extraction and LSTM for temporal beam prediction. | RMSE | |
+[6] | Conv-LSTM | The RSRP image of all wide beams | The RSRP image of all narrow beams | Using sub-pixel convolution and Conv-LSTM to realize RSRP prediction. | MSE | |
+Our | CS+AI | Channel information of compressed multi-beam | RSRP of all beams at the current instance | Learning the optimal measurement matrix and conv-LSTM network for prediction.
+| Top-1 accuracy | |
+[3] | Cell-level | CNN+LSTM | SNR of all beams and all cells | The same as the input | Using CNN to reduce dimension and LSTM network to predict beam quality. | link(s) prediction error |
+[15] | CNN/LSTM | RSRP images of all beams of all cells | The best BS index, and path state (LoS or NLoS) | Using two AI models based on CNN and LSTM predicts the best BS index.
+| Prediction accuracy of best BS | |
+Our | CNN/LSTM | Compressed RSRP images | RSRP of all beams of all cells | Using two AI models to optimize the shortcomings of the handover mechanism. | RLF rates |
+III-B AI-Based Proactive Handover Scheme
+Existing AI/ML-assisted cell handover methods typically use the RSRP of all beams across cells to predict the future RSRP for each beam or to identify the optimal base station index for upcoming handovers [15, 3].
+These methods primarily focus on the improvements in prediction accuracy and spectral efficiency provided by AI models, with limited attention to system-level metrics, such as RLF rates and PP rates.
+This gap may overlook the potential impact of model design on system reliability and service continuity.
+Additionally, existing algorithms do not sufficiently address the inherent limitations of the event A3-based handover mechanism.
+For instance, event A3-based handover is triggered and executed based on historical measurement reports, making it a passively reactive approach.
+The system initiates the handover only after the UE detects communication quality degradation, where the communication quality has already suffered irreversible loss.
+To maintain a stable communication quality for HSR communications, we provide a novel AI-assisted cell handover scheme.
+We use an AI model to predict the L3-RSRP of all cells for future state 1 based on historical L1-RSRP to aid the source BS in handover decisions.
+Using this method, we predict the future communication quality variation trends of adjacent cells and proactively switch cells based on these predictions.
+This approach effectively mitigates the communication quality degradation and reduces handover time from the original states (state 1, state 2, state 3) to the existing states (state 2, state 3).
+The AI-based cell handover process is illustrated in Fig.
+5, where option 1 and option 2 represent different deployment methods of AI models.
+In option 1, the AI-based prediction model is deployed on the UE side, with the prediction report uploaded to the source BS to assist in the handover decision.
+In option 2, the AI-based prediction model is deployed on the network side, where the network performs temporal prediction and makes the handover decision based on historical measurement reports.
+The primary difference between the two options lies in the model’s deployment mode, while the model’s inputs, outputs, and architecture remain the same.
+Based on the handover scheme we propose, we identify two approaches to reduce measurement overhead in cell-level mobility management.
+For the first approach, all beams from the partially selected cells are used to predict the L3-RSRP for all cells.
+For the second approach, the partially selected beams of each cell are used to predict the L3-RSRP for all cells.
+To be noticed, the model input is obtained through equal-interval sampling rather than the compressed multi-beam measurement method proposed in Section II to independently evaluate the performance gains of the AI-assisted handover scheme.
 We utilize the partial L1-RSRP from the past six time instances to predict the L3-RSRP in the next four time instances.
-We employ CNN-based and LSTM-based models as shown in Fig.
-5.
-The LSTM-based model comprises four LSTM layers and each layer has 168 hidden units.
-The features extracted by the LSTM network undergo further processing via an FC layer and a sigmoid function for feature mapping.
+We employ CNN-based and LSTM-based models for RSRP prediction.
+The LSTM-based model comprises four LSTM layers and two FC layers.
 To enhance the performance of the LSTM network, we adopt four identical networks to predict L3-RSRP for four future instances separately.
@@
 To avoid overfitting during the training process, max pooling layers are incorporated after each CNN layer.
-The channel used in our experiments is generated by QuaDRiGa, where we consider the UMa scenario and center frequency 30 GHz.
-We deploy 7 BSs, and each BS is equipped with three half-wavelength spaced uniform plane arrays (UPAs) with sector orientation of 30, 150 and -90 degrees.
+The channel used in our experiments is generated by QuaDRiGa, with the UMa scenario and a center frequency of 30 GHz, consistent with the beam-level experiment setup.
+We deploy 7 BSs, and each BS is equipped with three half-wavelength spaced uniform plane arrays (UPAs) with sector orientation of 30, 150 and -90 degrees and interference is generated by 21 adjacent cells, modeled as random interference.
 We generate four datasets corresponding to various UE speeds of 60 km/h, 120 km/h, 350 km/h, and 500 km/h.
@@
 In addition, we also consider the RLF rates as the system-level evaluation metric to illustrate the overall handover performance of the communication system.
-The RSRP difference between the AI predictions and the actual measurements is shown in Fig.
-6.
 The experimental results show that both CNN-based and LSTM-based models show excellent cell-level temporal prediction performance.
@@
 Remarkably, the proposed two approaches to reduce the measurement overhead for cell-level beam management, labeled as Part_Cell and Part_Beam, achieve comparable RLF performance with the traditional approach using all beam measurements of all cells, labeled as All_Beam_Cell, while we can save 50% beam measurement overhead for HSR communications.
-Furthermore, as shown in Table II, the flops and total parameters of the CNN-based model are much lower than those of the LSTM-based model.
+Furthermore, as shown in Table II, the flops and total parameters of the CNN-based model are much lower than those of the LSTM-based model, while achieving comparable performance.
 UE Speed | Flops | Total Params | ||||
@@
 Relying solely on RSRP variations is insufficient for ensuring communication stability during handover decisions.
-Therefore, a multi-criteria decision-making mechanism is imperative.
-This mechanism could integrate factors, such as signal interference levels, network load conditions, terminal velocity, and direction of movement, along with RSRP variations as inputs for AI models to obtain better handover decisions.
+Therefore, a multi-criterion decision-making mechanism becomes imperative.
+This mechanism could integrate factors, such as signal interference levels, network load conditions, terminal velocity, and direction of movement, along with RSRP variations as inputs to the AI models to obtain better handover decisions.
 This would greatly improve the accuracy and reliability, ensuring seamless communication for high-speed trains.
@@
 In this case, the UE reports measurement data to the source BS, which feeds the measurements into the AI/ML models for mobility management.
-Moreover, the transmission entails extra latency while the unavoidable transmission errors, occurring in reporting these measurements, affect the performance of AI models.
+Moreover, the transmission entails extra latency while the unavoidable transmission errors, occurring during the reporting of these measurements, affect the performance of AI models.
 Then, it would be beneficial to deploy AI models on the UE side.
-We can consider using CS or training lightweight AI models to reduce the computational complexity and hardware requirements for UEs.
+We can consider employing advanced AI models, such as lightweight AI models or transformer-based architectures, to enhance the computational efficiency and reduce the hardware requirements for the UEs.
 IV-D Model Generalization Design
@@
 Therefore, one challenge for AI-based mobility management for HSR communications is the design of AI models with good scenario generalization properties.
-Generally, there exists a trade-off between model generalization and model performance.
-Improving the generalization of the model, e.g., training the model with a mixed datasets from multiple scenarios, will introduce difficulties for the training of AI/ML models, as more complex and diverse knowledge are embedded in the data.
-As a result, the balance between model generalization and prediction accuracy according to the actual application scenario needs to be thoroughly examined in future discussions.
+Continuous learning is particularly beneficial in dynamic environments, enabling models to adapt to shifting channel conditions and movement patterns in real time.
+However, enhancing the model generalization through continuous learning incurs significant deployment overhead.
+Balancing deployment overhead and generalization remains a critical research in HSR communication.
 V Conclusion
-In this article, we investigated the fundamental issue of AI-based beam-level and cell-level mobility management for HSR communications, including the use cases, inputs, outputs and KPIs of AI models.
+In this article, we investigated the fundamental issue of AI-based beam-level and cell-level mobility management for HSR communications of AI models.
 For beam-level mobility management in HSR communications, a hybrid approach, combining AI and CS, demonstrated superior performance compared to conventional AI-based methods with the same measurement overhead.
 For cell-level mobility management in HSR communications, we unveiled the significant gain of AI-based methods for cell handover, while the measurement overhead can be reduced by using partial beam measurements, without compromise of the RLF rates.
-At last, the opportunities and challenges on AI-based mobility management are briefly discussed for further studies.
+At last, the opportunities and challenges on AI-based mobility management were briefly discussed for further studies.